OpenStudy (aaronandyson):

Two natural numbers differ by 3 .Find the numbers,if the sum of their reciprocal is 7/10 @sepeario

OpenStudy (welshfella):

1/x +1/y = 7/10 y - x = 3

OpenStudy (aaronandyson):

?????

OpenStudy (welshfella):

if we let the 2 numbers be x and y the above system of equations will hold. So you need to solve for x and y

OpenStudy (aaronandyson):

Why not x and x+3?

OpenStudy (welshfella):

oops - yes tha'ts easier! maybe i'm not too sharp today lol

OpenStudy (welshfella):

so you need to solve: